# Scaling Laws for Optimal Data Mixtures

### v1

@@ 3.4 Implementation details
In order to scale models, we change the hidden dimension size in the transformers $d$, keeping a fixed number of layers. To reduce the experimental cost, most of the experiments are done with a constant learning rate scheduler. This allows us to collect many points with varying numbers of tokens $D$ for each run, instead of one per run, which means that we can run more experiments and explore the space of domain weights more thoroughly. We also validate our findings when using cosine learning scheduler in Sec. 6, where we show that the scaling laws also extrapolate from small-scale to large-scale behavior in that case.
@@ 4 Predicting large-scale performance from small-scale experiments
Figure 3: Observed vs predicted loss for LLM pretraining on domains from the slimpajama dataset, NMM pretraining with multimodal domains, and LVM pretraining with image-caption domains. The scaling laws are fitted on small-scale models (blue points in the figure) and extrapolated to larger models. We display here the average loss over all domains for each modality. The MREfor each domain is reported in Tab. 2.
In this section, we demonstrate that (a) our scaling laws accurately capture the training data, and (b) generalize effectively to larger scales with significantly increased values of N and D. To this end, we fit the laws using small models trained with a small number of tokens, and we validate them on larger models with a large number of tokens.
We experiment with LLMs, trained with mixture of text domains, NMMs, trained with mixture of multimodal domains, and LVMs, trained with mixture of images and paired image-text domains. For LLMs, we consider the $k=7$ domains from slimpajama, which are different text domains. For multimodal pretraining, and similar to previous works, [34, 69, 55], the data mixture spans $k=3$ different domains: text, paired (image-captions), and interleaved multimodal data. For large vision model pretraining, we use $k=4$ domains.
Tab. 2 displays the different model sizes, number of training tokens, and number of different domains weights that we use to train and evaluate the scaling laws.
Results. Fig. 3 presents a comparison between the actual loss achieved by our trained models and the loss predicted by our scaling laws. We summarize the results on each domain by showing the average predicted loss (full results in appendix B). Remarkably, the predicted losses align closely with the observed values for both the joint and additive laws. In addition, the laws show good extrapolation to larger model sizes. To further quantify this alignment, we report the mean relative error (MRE%) in Tab. 2, which reveals a consistently low MREfor both laws, with an improvement of the joint law over the additive one. These results demonstrate that we can fit the scaling laws on small scales and extrapolate to larger scales. We remark that the MREs have some degree of variability across domains; for instance, on the LLM experiment, we get at the same time an extremely low MRE of $0.31\%$ on the C4 dataset and a high MRE of $4.45\%$ on Wikipedia.
@@
N D ##\## domain weights
LLM Train 412M 4-20B 60
@@
Eval 3B 20-100B 4
NMM Train 106M 20-100B 32
@@
Eval 2.3B 100-160B 8
8B 100-160B 4
LVM Train 89M 17-50B 32
@@
Modality Target domain MRE(Additive / Joint law)
@@ 5 Optimal data mixtures
$h^{*}(N,D)\in\arg\min_{h\in\Delta_{k}}\sum_{i=1}^{m}\mathcal{L}^{i}(N,D,h).$ (8)
The behavior of $h^{*}$ depends on the scaling law that we consider. Since eq. 4 assumes an additive relationship, the minimizer of eq. 8 is independent of $N,D$; in other words, it does not depend on scale. On the other hand, eq. 5 takes into account a multiplicative interaction between $N,D$ and $h$. Therefore, the optimal $h$ is scale-dependent. If one task is more important than another, we can incorporate importance weights in the sum in eq. 8.
The main practical takeaway of our paper is this simplified approach to optimal mixture estimation. Indeed, as demonstrated in Fig. 3, we can accurately fit our scaling laws with small-scale runs. Using these scaling laws, we can then solve eq. 8 for various targets $(N,D)$, which gives a principled way of choosing domain weights, rather than using ad-hoc methods as usually done in practice. To demonstrate our point, we do this for different modalities considered in the paper
Figure 4: Losses of the 7B models. After fitting the scaling laws on the small scale runs, we estimate the optimal domain weights $h^{*}_{avg}$ that minimize the average loss over the training domains (left), and $h^{*}_{OH}$ that minimizes the loss on the OpenHermes dataset (right). We then train 7B models with these optimal weights, and compare them to two baselines: one with uniform weights, and one with the standard distribution of slimpajama. The losses are averaged over all training domains, and also reported on the OpenHermes dataset. As expected, the model trained with $h^{*}_{OH}$ performs best on OpenHermes, while the model trained with $h^{*}_{avg}$ performs best on the training domains. Note that in this case $h^{*}_{avg}$ is close to uniform (Section D.1).
@@
Table 3: Evaluations of the 7B models. We report the median score on the CORE tasks [35] as well as MMLU. The model trained with weights that minimize the loss on OpenHermes performs best. Results on other benchmarks are in Appendix C.
Mixture weights CORE MMLU
Base 52 25
Uniform 53 30
$h^{*}_{avg}$ 56 32
$h^{*}_{OH}$ 58 37
For all the runs, we also monitor the loss on the OpenHermes dataset, which is a small high-quality dataset used for model alignment. We fit the scaling laws for that domain as well, even though this domain is not part of the pre-training domains. The rationale is that we want to estimate weights that lead to the best performance on this high quality dataset, which should be a proxy of performance on downstream tasks. We then find the optimal domain weights for that scaling law, which we denote $h^{*}_{OH}$, and train another 7B model with 150B tokens. As baselines, we train two more 7B models with as many tokens, one with the standard distribution of slimpajama, where domain weights are proportional to the number of tokens in each domain, and one with a uniform domain weights. Since we want the best models possible, we use a cosine learning rate schedule. We report the average loss on the training domains, and on the OpenHermes dataset in Figure 4. We also evaluate them on several downstream tasks and report the results in Table 3. The model trained with weights $h^{*}_{OH}$ is overall better than the other models. We believe that the pipeline demonstrated in this paper — estimate the scaling law for the loss on a high quality domain with small scale runs, find the minimizer, train a large scale model with it — is a promising avenue to get better models.
NMM results. We fit both the additive and joint scaling laws on three multimodal data domains, using only small models. For the joint scaling laws, we predict the best training mixture $h^{*}$ that minimizes the average of the domain losses for each model size while fixing the number of tokens at 100B for practicality. We then train models with these optimized mixtures. Fig. 1 compares the performance of these best mixtures against uniform mixtures, those used in prior works [55, 43], and randomly sampled mixtures that cover an important area of mixture grid. Models trained with our estimated mixtures consistently outperform alternatives. Notably, both additive and joint laws perform similarly well, making additive laws a strong and more practical baseline, since it takes the same optimal mixture for all runs. Remarkably, the optimized mixtures generalize effectively to larger model sizes, which validates the possibility of choosing the optimal mixture based on small-scale experiments, and then extrapolating to larger scales.
@@ 6 Scaling laws analysis
In that section, we conduct LLM experiments in a different setting, using the Pile dataset [21], which allows us to have a variable number of domains, between $k=4$ and $k=8$.
@@
We investigate the number of runs needed for an accurate scaling law fitting. We randomly partition the domain weights into $h_{\mathrm{train}}=[h_{1},\dots,h_{q}]$, of size $q$, and put the other domain weights into $h_{\mathrm{test}}$. We fit the scaling law on $h_{\mathrm{train}}$ and report the MRE on the test domain weights.
Since the number of parameters of the scaling law depends linearly on the number of domains $k$, we expect the number of runs necessary to fit the scaling law to increase when we consider more domains. To verify this hypothesis, we consider the NMM pretraining experiments, with $k=3$ domains and LLM pretraining with $k=4,6,8$ domains. For the NMM with 3 domains and LLM with 4 domains as considered so far, we fit the laws on small-scale models and compute the MRE on large-scale models as in Fig. 3. For the LLM training with $k=6,8$ domains, because of the very large search space with a high number of domains, we take a single model size and skip the dependency on $N$ in the scaling law, only considering the dependency on $h$ and the number of training tokens $D$. We report the MRE as a function of the number of training histograms $q$ in Fig. 5. We observe that we need about 10 runs for the NMM and LLM with 4 domains to get to an optimal MRE, while we need about 20 for LLM with 6 and 8 domains. Interestingly, we observe that when the number of training runs is very low, the additive law has a slightly lower eval MRE, due to its lower number of parameters.
@@
We study how the optimal mixture $h^{*}(N,D)$ for the average loss evolves as a function of the compute-budget $(N,D)$ on the multimodal models, as predicted by the joint scaling law. We report results in  Fig. 6. We see that interleaved data gets less important as we increase $D$, whereas bigger models tend to rely more on text. The additive law captures the average behavior across all scales.
Figure 6: Evolution of optimal domain weights $h^{*}$ with compute budget $(N,D)$ on the multimodal data, as predicted by the joint scaling law (eq. 5).
@@
The bulk of our experiments use a constant learning rate, which helps us gather many $D$ points for each run, but this departs from what is done in practice when training competitive models, where a cosine learning rate is typically used. In order to validate that our scaling laws are still valid when training with a cosine learning rate, we repeat the LLM experiments with $k=4$ pile domains with cosine learning rate decays, with fewer runs, training for $5$ different values of $D$, with $25$ different domain weights. We use $90M,200M,350M$ and $700M$ models for training, and extrapolate to $1.3B$. We observe that our scaling law fit is similar to those in the main experiments: on the 1B model, we get to an average MRE of $0.76\%$ for the additive law and $0.54\%$ for the joint law. We report detailed results in appendix B. Interestingly, the estimated optimal domain weights for the average loss are very similar to those estimated with the constant learning rate: for the additive law, we have $h^{*}_{\mathrm{cos}}=[0.35,0.18,0.30,0.17]$ and $h^{*}_{\mathrm{const}}=[0.34,0.17,0.32,0.17]$.
Other scaling laws formulas.
We investigate alternative scaling laws, and validate our proposed scaling laws, by evaluating other formulas in the same setup as in Fig. 3. First, we want to understand whether we could use a simpler form for the dependency on the domain weights $h$. To do so, we use the “simple additive” scaling law
$\mathcal{L}=E+(\sum_{i=1}^{k}C_{i}{h_{i}})^{\gamma}+\frac{A}{{D}^{\alpha}}+\frac{B}{{N}^{\beta}},$ (9)
where the dependency in $h$ is simpler compared to the additive and joint scaling laws. This law has $k-1$ fewer parameters than the additive scaling law. The joint scaling law models the terms $A^{h}$ and $B^{h}$ as functions of domain weights. We want to understand whether also taking a dependency of $\alpha$ and $\beta$ on $h$ helps capture more information about models’ behavior. To do so, we consider the “full” scaling law:
$\mathcal{L}$ $=E+\frac{1}{\sum_{i=1}^{k}C_{i}h_{i}^{\gamma_{i}}}+\frac{A^{h}}{N^{\alpha^{h}}}+\frac{B^{h}}{D^{\beta^{h}}}\text{, with }$ (10)
$A^{h}=(\sum_{i=1}^{k}C^{A}_{i}h_{i})^{\gamma^{A}},\enspace B^{h}=$ $(\sum_{i=1}^{k}C_{i}^{B}h_{i})^{\gamma^{B}},\enspace\alpha^{h}=(\sum_{i=1}^{k}C^{\alpha}_{i}h_{i})^{\gamma^{\alpha}}\text{ and }\beta^{h}=(\sum_{i=1}^{k}C^{\beta}_{i}h_{i})^{\gamma^{\beta}}$ (11)
This law is more expressive than the joint scaling law, and it adds $12k+$ 1 parameters. We give the full results of those laws in appendix B, and report the average MRE in Tab. 4. Overall, while the full law gives the best results for the LLM and LVM experiments, we believe that the simplicity of the additive and joint laws, combined with their relatively low MREs, makes them candidates of choice for scaling law formulas. For the LVM experiment, the picture is different: the simple law works better than the additive law, and the full scaling law brings a sizeable improvement over the joint law.
@@
$CE(p,q(h))$ $=\underbrace{H(p)+KL(p\|q(h^{*}))}_{\text{constant, independant of h}}$ $+$ $\underbrace{KL(p\|q(h))-KL(p\|q(h^{*}))}_{\geq 0\text{ by hypothesis on }h^{*}}$ (12)
@@
We can identify both terms since they are of the form “constant plus non-negative function that cancels”. We see that $E+(\sum_{i=1}^{k}C_{i}{h^{*}_{i}}^{\gamma_{i}})^{-1}$ captures both the intrinsic entropy $H(p)$ of the target distribution, and the shift $KL(p\|q(h^{*}))$ induced by training on the optimal mixture $h^{*}$, while the right-hand term is the expected log-likelihood ratio $\mathbb{E}_{p}[\log{\frac{q(h^{*})}{q(h)}}]$, which measures how far the model trained on $h$ is from the optimal one. If $p$ is one of the training domains $\mathcal{D}_{i}$, for disjoint domains we can assume $h^{*}_{i}\approx 1$ (see appendix D for justification) and simply bound its entropy $H(p)\leq E+C_{i}^{-1}$.

### v2

@@ 3.4 Implementation details
In order to scale models, we change the hidden dimension size in the transformers $d$, keeping a fixed number of layers. To reduce the experimental cost, most of the experiments are done with a constant learning rate scheduler. This allows us to collect many points with varying numbers of tokens $D$ for each run, instead of one per run, which means that we can run more experiments and explore the space of domain weights more thoroughly. We also validate our findings when using cosine learning scheduler in section˜6, where we show that the scaling laws also extrapolate from small-scale to large-scale behavior in that case.
@@ 4 Predicting large-scale performance from small-scale experiments
Figure 3: Observed vs predicted loss for LLM pretraining on domains from the slimpajama dataset, NMM pretraining with multimodal domains, and LVM pretraining with image-caption domains. The scaling laws are fitted on small-scale models (blue points in the figure) and extrapolated to larger models. We display here the average loss over all domains for each modality. The MRE% for each domain is reported in table˜2.
In this section, we demonstrate that (a) our scaling laws accurately capture the training data, and (b) generalize effectively to larger scales with significantly increased values of N and D. To this end, we fit the laws using small models trained with a small number of tokens, and we validate them on larger models with a large number of tokens. We experiment with LLMs, trained with a mixture of text domains, NMMs, trained with a mixture of multimodal domains, and LVMs, trained with images of different qualities and paired or not with text. For LLMs, we consider the $k=7$ domains from slimpajama, which are different text domains. For multimodal pretraining, and similar to previous works, (Laurençon et al., 2024; Zhang et al., 2024b; Shukor et al., 2025), the data mixture spans $k=3$ different domains: text, paired (image-captions), and interleaved multimodal data. For large vision model pretraining, we use $k=4$ domains.
table˜2 displays the different model sizes, number of training tokens, and number of different domains weights that we use to train and evaluate the scaling laws.
Results. figure˜3 presents a comparison between the actual loss achieved by our trained models and the loss predicted by our scaling laws. We summarize the results on each domain by showing the average predicted loss (full results in appendix˜B). Remarkably, the predicted losses align closely with the observed values for both the joint and additive laws. In addition, the laws show good extrapolation to larger model sizes. To further quantify this alignment, we report the mean relative error (MRE%) in table˜2, which reveals a consistently low MRE% for both laws, with an improvement of the joint law over the additive one. These results demonstrate that we can fit the scaling laws on small scales and extrapolate to larger scales. We remark that the MREs have some degree of variability across domains; for instance, on the LLM experiment, we get at the same time an extremely low MRE of $0.31\%$ on the C4 dataset and a high MRE of $4.45\%$ on Wikipedia.
FLOPs count. We report the approximate computational cost of running the small-scale experiments to fit the scaling laws, and contrast it with the cost of large-scale runs in table˜2. We compute the FLOPs required for one run as $6ND$. We observe that the cost of large-scale runs is comparable to that of small-scale runs. We also note that our large-scale runs could be trained with far more tokens, which would increase their FLOPs.
@@
N D #\# domain weights FLOPs
LLM Train 412M 4-20B 60 $2.5\times 10^{22}$
@@
Eval 3B 20-100B 4 $7.2\times 10^{21}$
7B 150B 1 $6.3\times 10^{21}$
NMM Train 106M 20-100B 32 $3.7\times 10^{22}$
@@
Eval 2.3B 100-160B 8 $1.8\times 10^{22}$
8B 100-160B 4 $3.1\times 10^{22}$
LVM Train 89M 17-50B 32 $1.4\times 10^{22}$
@@
Modality Target domain MRE% (Additive / Joint law)
@@ 5 Optimal data mixtures
$h^{*}(N,D)\in\arg\min_{h\in\Delta_{k}}\sum_{i=1}^{m}\mathcal{L}^{i}(N,D,h).$ (5.2)
The behavior of $h^{*}$ depends on the scaling law that we consider. Since equation˜2.4 assumes an additive relationship, the minimizer of equation˜5.2 is independent of $N,D$; in other words, it does not depend on scale. On the other hand, equation˜2.5 takes into account a multiplicative interaction between $N,D$ and $h$. Therefore, the optimal $h$ is scale-dependent. If one task is more important than another, we can incorporate importance weights in the sum in equation˜5.2.
The main practical takeaway of our paper is this simplified approach to optimal mixture estimation. Indeed, as demonstrated in figure˜3, we can accurately fit our scaling laws with small-scale runs. Using these scaling laws, we can then solve equation˜5.2 for various targets $(N,D)$, which gives a principled way of choosing domain weights, rather than using ad-hoc methods as usually done in practice. To demonstrate our point, we do this for different modalities considered in the paper
Figure 4: Losses of the 7B models. After fitting the scaling laws on the small scale runs, we estimate the optimal domain weights $h^{*}_{avg}$ that minimize the average loss over the training domains (left), and $h^{*}_{OH}$ that minimizes the loss on the OpenHermes dataset (right). We then train 7B models with these optimal weights, and compare them to two baselines: one with uniform weights, and one with the standard distribution of slimpajama. The losses are averaged over all training domains, and also reported on the OpenHermes dataset. As expected, the model trained with $h^{*}_{OH}$ performs best on OpenHermes, while the model trained with $h^{*}_{avg}$ performs best on the training domains.
@@
For all the runs, we also monitor the loss on the OpenHermes dataset, which is a small high-quality dataset used for model alignment. We fit the scaling laws for that domain as well, even though this domain is not part of the pre-training domains. The rationale is that we want to estimate weights that lead to the best performance on this high quality dataset, which should be a proxy of performance on downstream tasks. We then find the optimal domain weights for that scaling law, which we denote $h^{*}_{OH}$, and train another 7B model with 150B tokens. As baselines, we train two more 7B models with that many tokens, one with the standard distribution of slimpajama, proportional to the number of tokens in each domain, and one with a uniform distribution over domains.
Since we want the best models possible, we use a cosine learning rate schedule, which makes the scaling law extrapolation impossible to conduct. We report the average loss of these models on the training domains, on the OpenHermes dataset in Figure˜4. We also evaluate them on many downstream tasks and report the results in Table˜3. The model trained with weights $h^{*}_{OH}$ is overal better than the other models in terms of evaluations. We believe that the pipeline demonstrated in this paper — estimate the scaling law for the loss on a high quality domain with small scale runs, find the minimizer, train a large scale model with it — is a promising avenue to get better models.
Table 3: Evaluations of the 7B models. We report the median score on the CORE tasks (Li et al., 2024) as well as other accuracies on standard benchmarks. The model trained with weights that attempt to minimize the loss on OpenHermes performs best.
$h^{*}_{avg}$ 56 32 71 40 60 79 52 72 65
$h^{*}_{OH}$ 58 37 70 40 67 79 53 72 65
Uniform 53 30 70 40 46 78 51 70 65
Base 52 25 70 43 51 79 53 71 65
NMM results. We fit both the additive and joint scaling laws on three multimodal data domains, using only small models. For the joint scaling laws, we predict the best training mixture $h^{*}$ that minimizes the average of the domain losses for each model size while fixing the number of tokens at 100B for practicality. We then train models with these optimized mixtures. figure˜1 compares the performance of these best mixtures against uniform mixtures, those used in prior works (Shukor et al., 2025; McKinzie et al., 2024), and randomly sampled mixtures that cover an important area of mixture grid. Models trained with our estimated mixtures consistently outperform alternatives. Notably, both additive and joint laws perform similarly well, making additive laws a strong and more practical baseline, since it takes the same optimal mixture for all runs. Remarkably, the optimized mixtures generalize effectively to larger model sizes, which validates the possibility of choosing the optimal mixture based on small-scale experiments, and then extrapolating to larger scales.
@@ 6 Scaling laws analysis
Figure 5: Evolution of optimal domain weights $h^{*}$ with compute budget $(N,D)$ on the multimodal data, as predicted by the joint scaling law (equation˜2.5).
@@
In that section, we conduct LLM experiments in a different setting, using the Pile dataset (Gao et al., 2020), which allows us to have a variable number of domains, between $k=4$ and $k=8$.
@@
We investigate the number of runs needed for an accurate scaling law fitting. We randomly partition the domain weights into $h_{\mathrm{train}}=[h_{1},\dots,h_{q}]$, of size $q$, and put the other domain weights into $h_{\mathrm{test}}$. We fit the scaling law on $h_{\mathrm{train}}$ and report the MRE on the test domain weights. Since the number of parameters of the scaling law depends linearly on the number of domains $k$, we expect the number of runs necessary to fit the scaling law to increase when we consider more domains. To verify this hypothesis, we consider the NMM pretraining experiments, with $k=3$ domains and LLM pretraining with $k=4,6,8$ domains. For the NMM with 3 domains and LLM with 4 domains as considered so far, we fit the laws on small-scale models and compute the MRE on large-scale models as in figure˜3. For the LLM training with $k=6,8$ domains, because of the very large search space with a high number of domains, we take a single model size and skip the dependency on $N$ in the scaling law, only considering the dependency on $h$ and the number of training tokens $D$. We report the MRE as a function of the number of training histograms $q$ in figure˜6. We observe that we need about 10 runs for the NMM and LLM with 4 domains to get to an optimal MRE, while we need about 20 for LLM with 6 and 8 domains. Interestingly, we observe that when the number of training runs is very low, the additive law has a slightly lower eval MRE, due to its lower number of parameters.
@@
We study how the optimal mixture $h^{*}(N,D)$ for the average loss evolves as a function of the compute-budget $(N,D)$ on the multimodal models, as predicted by the joint scaling law. We report results in figure˜5. We see that interleaved data gets less important as we increase $D$, whereas bigger models tend to rely more on text. The additive law captures the average behavior across all scales.
@@
The bulk of our experiments use a constant learning rate, which helps us gather many $D$ points for each run, but this departs from what is done in practice when training competitive models, where a cosine learning rate is typically used. In order to validate that our scaling laws are still valid when training with a cosine learning rate, we repeat the LLM experiments with $k=4$ pile domains with cosine learning rate decays, with fewer runs, training for $5$ different values of $D$, with $25$ different domain weights. We use $90M,200M,350M$ and $700M$ models for training, and extrapolate to $1.3B$. We observe that our scaling law fit is similar to those in the main experiments: on the 1B model, we get to an average MRE of $0.76\%$ for the additive law and $0.54\%$ for the joint law. We report detailed results in appendix˜B. Interestingly, the estimated optimal domain weights for the average loss are very similar to those estimated with the constant learning rate: for the additive law, we have $h^{*}_{\mathrm{cos}}=[0.35,0.18,0.30,0.17]$ and $h^{*}_{\mathrm{const}}=[0.34,0.17,0.32,0.17]$.
@@
Other scaling laws formulas.
We investigate alternative scaling laws, and validate our proposed scaling laws, by evaluating other formulas in the same setup as in figure˜3. First, we want to understand whether we could use a simpler form for the dependency on the domain weights $h$. To do so, we use the “simple additive” scaling law
$\mathcal{L}=E+(\sum_{i=1}^{k}C_{i}{h_{i}})^{\gamma}+\frac{A}{{D}^{\alpha}}+\frac{B}{{N}^{\beta}},$ (6.1)
where the dependency in $h$ is simpler compared to the additive and joint scaling laws. This law has k1k-1 fewer parameters than the additive scaling law. The joint scaling law models the terms $A^{h}$ and $B^{h}$ as functions of domain weights. We want to understand whether also taking a dependency of $\alpha$ and $\beta$ on $h$ helps capture more information about models’ behavior. To do so, we consider the “full” scaling law:
$\mathcal{L}$ $=E+\frac{1}{\sum_{i=1}^{k}C_{i}h_{i}^{\gamma_{i}}}+\frac{A^{h}}{N^{\alpha^{h}}}+\frac{B^{h}}{D^{\beta^{h}}}\text{, with }$ (6.2)
$A^{h}=(\sum_{i=1}^{k}C^{A}_{i}h_{i})^{\gamma^{A}},\kern 5.0ptB^{h}=$ $(\sum_{i=1}^{k}C_{i}^{B}h_{i})^{\gamma^{B}},\kern 5.0pt\alpha^{h}=(\sum_{i=1}^{k}C^{\alpha}_{i}h_{i})^{\gamma^{\alpha}}\text{ and }\beta^{h}=(\sum_{i=1}^{k}C^{\beta}_{i}h_{i})^{\gamma^{\beta}}$ (6.3)
This law is more expressive than the joint scaling law, and it adds $2k+1$ parameters. We give the full results of those laws in appendix˜B, and report the average MRE in table˜4. We see that the additive law reduces the MRE a lot compared to the simple law, especially in the LLM experiment. Despite having a slightly smaller train MRE, the full law does not extrapolate as well as the joint law. For the LVM experiment, the picture is different: the simple law works better than the additive law, and the full scaling law brings some improvement over the joint law.
@@
$CE(p,q(h))$ $=\underbrace{H(p)+KL(p\|q(h^{*}))}_{\text{constant, independant of h}}$ $+$ $\underbrace{KL(p\|q(h))-KL(p\|q(h^{*}))}_{\geq 0\text{ by hypothesis on }h^{*}}$ (6.4)
@@
We can identify both terms since they are of the form “constant plus non-negative function that cancels”. We see that $E+(\sum_{i=1}^{k}C_{i}{h^{*}_{i}}^{\gamma_{i}})^{-1}$ captures both the intrinsic entropy $H(p)$ of the target distribution, and the shift $KL(p\|q(h^{*}))$ induced by training on the optimal mixture $h^{*}$, while the right-hand term is the expected log-likelihood ratio $\mathbb{E}_{p}[\log{\frac{q(h^{*})}{q(h)}}]$, which measures how far the model trained on $h$ is from the optimal one. If $p$ is one of the training domains $\mathcal{D}_{i}$, for disjoint domains we can assume $h^{*}_{i}\approx 1$ (see appendix˜C for justification) and simply bound its entropy $H(p)\leq E+C_{i}^{-1}$.